A coffee shop mixes French roast coffee worth $9 per pound with Kona coffee worth $4 per pound. How much of each type of coffee was used in order to obtain 20 pounds of a mixture worth $5 per pound. Thank You
Answer by checkley77(12844)

9*x+4(20-x)=5*20
9x+80-4x=100
5x=100-80
5x=20
x=20/5
x=4 pounds of $9 coffee is used.
20-4=16 pounds of $4 coffee is used.
Proof:
9*4+4*16=5*20
36+64=100
100=100